OBSERVATION OF A MODERATE STRENGTH INTERACTION OF HYDROGEN WITH A COINAGE METAL HALIDE: THE ROTATIONAL SPECTRUM AND STRUCTURE OF THE p-H_{2}-CuCl AND o-H_{2}- CuCl COMPLEXES

\underline{HERBERT M. PICKETT}, DANIEL A. OBENCHAIN, G. S. GRUBBS II, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, 52 Lawn Avenue, Middletown, CT, 06459-0180, USA.

Rotational transitions of the p-H_{2}-CuCl and o-H_{2}-CuCl have been observed on a laser ablation equipped FTMW cavity instrument. Computational studies preformed using the APFD^{a} density functional and MP2 level of theory were used to predict the structure of the p-H_{2}-CuCl. Measurements from the J=1–0 to the J=3–2 transitions were used to determine the rotational constants, centrifugal distortion constants, and quadrupole coupling constants for multiple isotopologues of the p-H_{2}-CuCl species. Similar constants, including spin-spin coupling constants, have also been determined for the o-H_{2}-CuCl species for the J=2–1 and the J=3–2 transitions. The eQq of the copper in p-H_{2}-^{63}Cu^{35}Cl was found to be 52.058(2) MHz, a change from the monomer ^{63}Cu^{35}Cl value of 16.1712(24) MHz^{b}.

aA. Austin, G. A. Petersson, M. J. Frisch, F. J. Dobek, G. Scalmani, and K. J. Throssell. Chem. Theor. Comp. 8 (2012) 4989.

bK. D. Hensel, C. Styger, W. Jager, A. J. Merer, and M. C. L. Gerry, J. Chem. Phys. 99(1993) 3320.